# Regularity of vector fields with piecewise regular curl and divergence

### v2

@@ 4.3 Helmholtz decompositions with piecewise regularity properties
Moreover, ${\bf v}:={\bf u}-\boldsymbol{\nu}^{-1}\operatorname{curl}{\bf z}$ satisfies $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=\operatorname{div}\boldsymbol{\nu}{\bf u}$, thus according to (i) there holds $\left\|{\bf v}\right\|_{{\bf PH}^{1}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)}$. If $\boldsymbol{\nu}$ is continuously differentiable on $\overline{\Omega}$, then (ii) is proved analogously, the only change is that the equations $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=\operatorname{div}\boldsymbol{\nu}{\bf u}$ then imply $\left\|{\bf v}\right\|_{{\bf H}^{1}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)}$. ∎
@@
Let $\Omega\subseteq\mathbb{R}^{3}$ be a bounded Lipschitz domain. There exist pseudodifferential operators $\operatorname{R}_{1}^{\Omega}$, ${\bf R}_{2}^{\Omega}$ of order $-1$ and ${\bf K}^{\Omega}$, ${\bf K}_{2}^{\Omega}$ of order $\infty$- ∞ on $\mathbb{R}^{3}$ which for $m\in\mathbb{Z}$ have the mapping properties
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},\quad\left\|\varphi\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
We notice that (5.4) implies $-\operatorname{div}\boldsymbol{\nu}\nabla\psi=\operatorname{div}(\boldsymbol{\nu}{\bf z})$ on every subdomain $\mathcal{G}_{i}$, and due to ${\bf v}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=0$ we have $\llbracket\nabla\psi\rrbracket^{\operatorname{t}}=-\llbracket{\bf z}\rrbracket^{\operatorname{t}}$ and $\llbracket\boldsymbol{\nu}\nabla\psi\rrbracket^{\operatorname{n}}=-\llbracket\boldsymbol{\nu}{\bf z}\rrbracket^{\operatorname{n}}$ on $\mathcal{I}$. Furthermore, the boundary condition $\boldsymbol{\nu}{\bf v}\cdot{\bf n}=0$ leads to $\boldsymbol{\nu}\nabla\psi\cdot{\bf n}=-\boldsymbol{\nu}{\bf z}\cdot{\bf n}$ on $\Gamma$.
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{1}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)},\quad\left\|{\bf z}\right\|_{{\bf H}^{1}(\Omega)}\leq\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
$\left\|{\bf r}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},\quad\left\|\varphi\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
hence due to $-k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ we get
@@
$\left\|{\bf u}_{T}\right\|_{{\bf H}^{m+1/2}(\Gamma)}\leq C\left(\left\|\operatorname{div}_{\Gamma}\boldsymbol{\zeta}{\bf u}_{T}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}+\left\|\operatorname{curl}_{\Gamma}{\bf u}_{T}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right),$
and due to $\operatorname{curl}_{\Gamma}{\bf u}_{T}=\operatorname{curl}{\bf u}\cdot{\bf n}$ on $\Gamma$, there holds $\left\|\operatorname{curl}_{\Gamma}{\bf u}_{T}\right\|_{{\bf H}^{m-1/2}(\Gamma)}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}$. In addition, taking the surface divergence on the impedance boundary condition yields
@@
$\left\|\operatorname{div}_{\Gamma}\boldsymbol{\zeta}{\bf u}_{T}\right\|_{{\bf H}^{m-1/2}(\Gamma)}\leq C\left(|k|\left\|{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+|k|^{-1}\left\|\operatorname{div}{\bf f}\right\|_{\operatorname{PH}^{m}(\mathcal{G})}+|k|^{-1}|\left\|\operatorname{div}_{\Gamma}{\bf g_{\operatorname{I}}}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right).$
@@
$\left\|{\bf u}_{T}\right\|_{{\bf H}^{m+1/2}(\Gamma)}\leq C\left(\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+|k|\left\|{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+|k|^{-1}\left\|\operatorname{div}{\bf f}\right\|_{\operatorname{PH}^{m}(\mathcal{G})}+|k|^{-1}\left\|\operatorname{div}_{\Gamma}{\bf g_{\operatorname{I}}}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right),$
@@
Case 2: Suppose that natural boundary conditions (1.1b) are imposed. By Theorem 2.5 and the equality $-k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ we have
@@
$\left\|\boldsymbol{\varepsilon}{\bf u}\cdot{\bf n}\right\|_{\operatorname{H}^{m+1/2}(\Gamma)}\leq|k|^{-2}\left\|{\bf f}\cdot{\bf n}\right\|_{\operatorname{H}^{m+1/2}(\Gamma)}+|k|^{-2}\left\|\operatorname{div}_{\Gamma}{\bf g_{\operatorname{N}}}\right\|_{\operatorname{H}^{m+1/2}(\Gamma)}.$ (6.4)
@@
Case 3: Suppose that essential boundary conditions (1.1c) are imposed. Due to Theorem 2.6 as well as $-k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ and ${\bf u}_{T}=0$ on $\Gamma$ we have
@@
$\vartheta_{\ell+1,k}\leq C_{\ell}\left(\vartheta_{0,k}+|k|^{-2}\sum_{m=0}^{\ell}\left[{\bf F}_{m,k}+{\bf G}_{m,k}\right]\right).$
@@
$\vartheta_{\ell+1,k}\leq C_{\ell}\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{-1}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{-2}\sum_{m=0}^{\ell}\left[{\bf F}_{m,k}+{\bf G}_{m,k}\right]\right).$
Multiplication by $|k|^{\ell+1}$ then concludes the proof of (2.1). Finally, in combination with (2.1), repeating Step 1 of the proof with $m=\ell+1$ yields (2.2) and thus finishes the proof of Theorem 2.7. ∎
@@ Acknowledgements
We gladly acknowledge financial support by Austrian Science Fund (FWF) throughout the doctoral school Dissipation and dispersion in nonlinear PDEs (grant W1245; MB) special research program Taming complexity in PDE systems (grant SFB F65, DOI:10.55776/F65; JMM).

### v3

@@ 4.3 Helmholtz decompositions with piecewise regularity properties
Moreover, ${\bf v}:={\bf u}-\boldsymbol{\nu}^{-1}\operatorname{curl}{\bf z}$ satisfies $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=\operatorname{div}\boldsymbol{\nu}{\bf u}$, thus according to (i) there holds $\left\|{\bf v}\right\|_{{\bf PH}^{1}{(\mathcal{G})}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)}$. If $\boldsymbol{\nu}$ is continuously differentiable on $\overline{\Omega}$, then (ii) is proved analogously, the only change is that the equations $\left({\bf v},\operatorname{curl}{\bf w}\right)_{{\bf L}^{2}(\Omega)}=0$ for all ${\bf w}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=\operatorname{div}\boldsymbol{\nu}{\bf u}$ then imply $\left\|{\bf v}\right\|_{{\bf H}^{1}{(\Omega)}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)}$. ∎
@@
Let $\Omega\subseteq\mathbb{R}^{3}$ be a bounded Lipschitz domain. There exist pseudodifferential operators $\operatorname{R}_{1}^{\Omega}$, ${\bf R}_{2}^{\Omega}$ of order 1-1 and ${\bf K}^{\Omega}$, ${\bf K}_{2}^{\Omega}$ of order -$\infty$ on $\mathbb{R}^{3}$ which for $m\in\mathbb{Z}$ have the mapping properties
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},\quad\left\|{\phi}\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
We notice that (5.4) implies $\operatorname{div}\boldsymbol{\nu}\nabla\psi=\operatorname{div}(\boldsymbol{\nu}{\bf z})$ on every subdomain $\mathcal{G}_{i}$, and due to ${\bf v}\in{\bf H}(\operatorname{curl},\Omega)$ and $\operatorname{div}\boldsymbol{\nu}{\bf v}=0$ we have $\llbracket\nabla\psi\rrbracket^{\operatorname{t}}=-\llbracket{\bf z}\rrbracket^{\operatorname{t}}$ and $\llbracket\boldsymbol{\nu}\nabla\psi\rrbracket^{\operatorname{n}}=-\llbracket\boldsymbol{\nu}{\bf z}\rrbracket^{\operatorname{n}}$ on $\mathcal{I}$. Furthermore, the boundary condition $\boldsymbol{\nu}{\bf v}\cdot{\bf n}=0$ leads to $\boldsymbol{\nu}\nabla\psi\cdot{\bf n}=-\boldsymbol{\nu}{\bf z}\cdot{\bf n}$ on $\Gamma$.
@@
$\left\|{\bf v}\right\|_{{\bf PH}^{1}{(\mathcal{G})}}\leq C\left\|\operatorname{div}\boldsymbol{\nu}{\bf u}\right\|_{\operatorname{L}^{2}(\Omega)},\quad\left\|{\bf z}\right\|_{{\bf H}^{1}(\Omega)}\leq\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
$\left\|{\bf r}\right\|_{{\bf PH}^{1}(\mathcal{G})}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)},\quad\left\|{\phi}\right\|_{\operatorname{H}^{1}(\Omega)}\leq C\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)},$
@@
hence due to $k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ we get
@@
$\left\|{\bf u}_{T}\right\|_{{\bf H}^{m+1/2}_{T}(\Gamma)}\leq C\left(\left\|\operatorname{div}_{\Gamma}\boldsymbol{\zeta}{\bf u}_{T}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}+\left\|\operatorname{curl}_{\Gamma}{\bf u}_{T}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right),$
and due to $\operatorname{curl}_{\Gamma}{\bf u}_{T}=\operatorname{curl}{\bf u}\cdot{\bf n}$ on $\Gamma$, there holds $\left\|\operatorname{curl}_{\Gamma}{\bf u}_{T}\right\|_{{\operatorname{H}^{m-1/2}(\Gamma)}}\leq C\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}$. In addition, taking the surface divergence on the impedance boundary condition yields
@@
$\left\|\operatorname{div}_{\Gamma}\boldsymbol{\zeta}{\bf u}_{T}\right\|_{{\bf H}^{m-1/2}(\Gamma)}\leq C\left(|k|\left\|{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+{|k|^{-1}\left\|{\bf f}\right\|_{{\bf PH}^{m}(\mathcal{G})}}+|k|^{-1}\left\|\operatorname{div}{\bf f}\right\|_{\operatorname{PH}^{m}(\mathcal{G})}+|k|^{-1}|\left\|\operatorname{div}_{\Gamma}{\bf g_{\operatorname{I}}}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right).$
@@
$\left\|{\bf u}_{T}\right\|_{{\bf H}^{m+1/2}(\Gamma)}\leq C$ $\left(\left\|\operatorname{curl}{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+|k|\left\|{\bf u}\right\|_{{\bf PH}^{m}(\mathcal{G})}+{|k|^{-1}\left\|{\bf f}\right\|_{{\bf PH}^{m}(\mathcal{G})}}+|k|^{-1}\left\|\operatorname{div}{\bf f}\right\|_{\operatorname{PH}^{m}(\mathcal{G})}\right$.
$\hskip 256.0748pt\left.+|k|^{-1}\left\|\operatorname{div}_{\Gamma}{\bf g_{\operatorname{I}}}\right\|_{\operatorname{H}^{m-1/2}(\Gamma)}\right),$
@@
Case 2: Suppose that natural boundary conditions (1.1b) are imposed. By Theorem 2.6 and the equality $k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ we have
@@
$\left\|\boldsymbol{\varepsilon}{\bf u}\cdot{\bf n}\right\|_{\operatorname{H}^{m+1/2}(\Gamma)}\leq|k|^{-2}\left\|{\bf f}\cdot{\bf n}{-\operatorname{div}_{\Gamma}{\bf g_{\operatorname{N}}}}\right\|_{\operatorname{H}^{m+1/2}(\Gamma)}.$ (6.4)
@@
Case 3: Suppose that essential boundary conditions (1.1c) are imposed. Due to Theorem 2.7 as well as $k^{2}\operatorname{div}\boldsymbol{\varepsilon}{\bf u}=\operatorname{div}{\bf f}$ and ${\bf u}_{T}=0$ on $\Gamma$ we have
@@
$\vartheta_{{m}+1,k}\leq C_{{m}}\left(\vartheta_{0,k}+|k|^{-2}\sum_{j=0}^{{m}}\left[{\bf F}_{j,k}+{\bf G}_{j,k}\right]\right).$
@@
$\vartheta_{{m}+1,k}\leq C_{{m}}\left(\left\|{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{-1}\left\|\operatorname{curl}{\bf u}\right\|_{{\bf L}^{2}(\Omega)}+|k|^{-2}\sum_{j=0}^{{m}}\left[{\bf F}_{j,k}+{\bf G}_{j,k}\right]\right)$
for all $m\in\{0,\ldots,\ell\}$. This is precisely (2.3), hence the proof of Theorem 2.10 is complete. ∎
@@ Acknowledgements
We gladly acknowledge financial support by Austrian Science Fund (FWF) throughout the special research program Taming complexity in PDE systems (grant SFB F65, DOI:10.55776/F65; JMM).